# Equivariant Symmetry-Aware Head Pose Estimation for Fetal MRI

### v4

@@ 4 Methods
Our goal is to estimate 6-DoF pose $T$ from a 3D volume, relative to a canonical object coordinate frame. As detailed in Fig. 2, we estimate pose by separating the rigid transform $T\!=\!t\circ R$, into its rotation $R$ and translation $t$ components.
@@ 4.1 Translation estimation
In rigid pose estimation, it is customary to place the origin of the object coordinate frame at the object center [xiang2018]. If $\psi:\mathbb{R}^{3}\rightarrow\{0,1\}$ is a binary object occupancy function, then $t$ can be computed from a set $X$ of sampled points (i.e., voxel coordinates):
$t=\frac{\sum_{x\in X}x\psi(x)}{\sum_{x\in X}\psi(x)}.$ (5)
See Appendix C for our implementation of $\psi$ using a network trained to segment the fetal brain segmentation from whole-uterus MRI volumes.
@@ 4.2 Rotation estimation
Let $G_{\text{symm}}\!=\!\{g_{s}\in\text{O(3)}\!:\!f\!\circ\!g_{s}^{-1}\!=\!f\}$ be the object symmetry group [saxena2007]. To handle object symmetries, we aim to parametrize $h(R)$ such that $\rho_{h}(g_{s})h(R)\!=\!h(R)$ for all $g_{s}\!\in\!G_{\text{symm}}$ and $R\!\in\!\text{SO(3)}$; and $h$ remains continuous over the quotient group $H\!=\!\text{O(3)}/G_{\text{symm}}$. Here, we consider the case $G_{\text{symm}}\!=\!\{g_{\leftrightarrow}\}$, where $g_{\leftrightarrow}$ denotes left-right reflection, defined by the axis of left-right symmetry $e_{\leftrightarrow}$. In this case, $H\!=\!\text{O(3)}/\{g_{\leftrightarrow}\}\!=\!\text{SO(3)}$.
Formally, we define $h(R)$ based on three orthonormal basis directions $\{e_{x},e_{y},e_{z}\}$ of the object frame (d=9d\!=\!9), a continuous function over SO(3):
@@
where $\oplus$ denotes columnwise and block-diagonal concatenation for vectors and matrices, respectively. Specifically, we decompose $h(R)$ into three irreducible tensors: we model the basis direction parallel to the left-right axis as a pseudovector, and the two basis directions orthogonal to the left-right axis as vectors. The invariance of $h(R)$ under $G_{\text{symm}}$ is straightforward:
$\rho_{h}(g_{\leftrightarrow})h(R)$ $=\!(\det{M(g_{\leftrightarrow})})M(g_{\leftrightarrow})e_{x}$ (8)
$\hskip 14.22636pt\!\oplus\!M(g_{\leftrightarrow})e_{y}\!\oplus\!M(g_{\leftrightarrow})e_{z},$
@@
where the second equality follows from the fact that vectors orthogonal and parallel to the axis of reflection remain unchanged and become inverted following the reflection, respectively. By using an E(3)-equivariant function to estimate $h(R)$ and further imposing invariance of $h(R)$ under $G_{\text{symm}}$, we ensure that $h(R)$ is equivariant under $\text{E(3)}/G_{\text{symm}}\!=\!\text{SE(3)}$ and continuous over SO(3).
@@
where $(e_{x},e_{y},e_{z})$ is the ground-truth (GT) orthonormal basis, $\theta_{k}\!=\!\arccos{(\hat{e}_{k}\cdot e_{k})}$ for $k\in\{x,y,z\}$, and $\beta$ is a weight hyperparameter. We use $\sin{(\theta/2)}$ for the standard vectors since it is monotonically related to the cosine similarity loss. We modify this term to $\sin{(\theta)}$ for the pseudovector in order to guarantee invariance under inversions (i.e., left-right reflections) of $e_{x}$. During inference, we obtain an orthonormal output matrix by applying SVD to the predicted basis [levinson2020]. We then choose between $\hat{e}_{x}$ and $\hat{e}_{x}$ in order to guarantee a positive determinant matrix, i.e., a proper rotation. This step is related to prior work that heuristically selects the output pose from a set of symmetrically plausible poses [pitteri2019, hodan2020, zhao2023]. Here, the task is straightforward since we are choosing between a right-handed and left-handed rotation.
@@ 4.3 Implementation details
We augment training volumes with commonly used spatial (rigid transformations, scaling, flips) and intensity (bias field, gamma, Gaussian noise, low resolution) transforms [perezgarcia2021, billot2023] in volumetric brain MRI. To enable our method to generalize to navigator volumes interleaved with diagnostic slices, we additionally simulate a spin history artifact from the preceding slice [miller2015, gagoski2016], which appears as dark shading along the slice imaging plane (blue arrows in Fig. 3). We model this artifact with a Gaussian approximation of the slice profile:
$\tilde{f}(x)$ $=f(x)(1-\mathcal{N}((x-c_{\text{slice}})^{T}n_{\text{slice}};0,\sigma^{2}),$ (10)
where $f$ and $\tilde{f}$ are the original and augmented volumes, respectively; $c_{\text{slice}}$ and $n_{\text{slice}}$ are the center and normal to the slice imaging plane, sampled uniformly at random; and $\sigma^{2}$ is determined by slice thickness and the rate of spin decay [kuklisovamurgasova2012] (see Appendix D for more details).
Table 1: Baseline comparisons. Mean $\pm$ standard deviation for rotation error () and average absolute error (AAD, $m$). Best score is in bold. * indicates statistical significance compared to E(3)-Pose (p<0.05p\!<\!0.05, pairwise Wilcoxon test). FireANTs and Fetal-Align rely only on segmentation networks trained on Research-Fetal and are not re-trained on dHCP, and therefore, we omit the corresponding rows for these methods. FireANTs runtime is 1.8s, prohibiting real-time deployment. EquiTrack runtime is 0.7s, and the remaining methods run in 0.3s.
@@
FireANTs [jena2024] 10.6 $\pm$ 24.5* 5.7 $\mathbf{\pm}$ 9.3* 0.4 $\mathbf{\pm}$ 1.6* 0.6 $\mathbf{\pm}$ 1.0* 17.3 $\pm$ 40.6 4.1 $\pm$ 7.4 44.6 $\pm$ 55.* 19.1 $\pm$ 20.3*
EquiTrack [billot2024] 11.3 $\pm$ 13.1* 6.2 $\pm$ 5.9* 7.1 $\pm$ 5.0 3.6 $\pm$ 2.6 18.9 $\pm$ 29.0* 5.4 $\pm$ 5.7* 44.7 $\pm$ 48.7* 20.3 $\pm$ 18.3*
Fetal-Align [hoffmann2021] 5.0 $\mathbf{\pm}$ 3.0 3.0 $\mathbf{\pm}$ 1.7 8.5 $\pm$ 24.7* 3.5 $\pm$ 6.8* 10.3 $\pm$ 21.0* 3.2 $\pm$ 5.0* 56.2 $\pm$ 57.6* 24.1 $\pm$ 23.3*
3DPose-Net [salehi2019] 34.0 $\pm$ 36.0* 16.9 $\pm$ 15.5* 29.5 $\pm$ 35.6* 13.0 $\pm$ 14.1* 35.7 $\pm$ 35.4* 10.2 $\pm$ 8.1* 48.1 $\pm$ 40.6* 22.1 $\pm$ 15.3*
6DRep [faghihpirayesh2023] 11.8 $\pm$ 6.8* 6.4 $\pm$ 3.5* 12.5 $\pm$ 7.1* 6.0 $\pm$ 3.4* 15.7 $\pm$ 9.1* 4.9 $\pm$ 2.6* 33.2 $\pm$ 9.5* 15.1 $\pm$ 13.8*
RbR [gopinath2024] 7.6 $\pm$ 3.9* 4.3 $\pm$ 2.2* 8.6 $\pm$ 5.0 4.3 $\pm$ 2.6 11.2 $\pm$ 4.6* 3.7 $\pm$ 1.4* 22.6 $\pm$ 24.1* 11.5 $\pm$ 9.9*
@@
EquiTrack [billot2024] 29.5 $\pm$ 43.5* 14.1 $\pm$ 19.5* 13.6 $\pm$ 36.7* 5.4 $\pm$ 11.6* 48.3 $\pm$ 62.1* 11.6 $\pm$ 13.2* 59.0 $\pm$ 57.5* 25.7 $\pm$ 22.1*
3DPose-Net [salehi2019] 64.1 $\pm$ 44.6* 29.9 $\pm$ 18.7* 38.4 $\pm$ 34.0* 17.0 $\pm$ 13.2* 74.9 $\pm$ 52.9* 18.9 $\pm$ 10.3* 87.3 $\pm$ 45.9* 36.1 $\pm$ 16.4*
6DRep [faghihpirayesh2023] 19.2 $\pm$ 9.9* 10.4 $\pm$ 5.5* 10.8 $\pm$ 5.5* 5.2 $\pm$ 2.6* 45.1 $\pm$ 44.9* 11.9 $\pm$ 9.8* 84.9 $\pm$ 57.5* 33.0 $\pm$ 17.4*
RbR [gopinath2024] 15.4 $\pm$ 18.3* 8.3 $\pm$ 8.5* 8.0 $\pm$ 4.0* 3.9 $\pm$ 1.9 50.9 $\pm$ 48.4* 12.7 $\pm$ 9.9* 77.2 $\pm$ 53.6* 30.8 $\pm$ 17.3*
@@ 5.1 Data
We train two variants of E(3)-Pose on two different large, high-quality research datasets. We first evaluate on in-distribution held-out test data. Then, we evaluate on two clinical datasets to assess domain generalization and the potential for clinical translation. All volumes are manually annotated for ground-truth pose and brain/eyes segmentations (see Appendix E for details).
@@
We train one instance of E(3)-Pose on Research-Fetal, a dataset of 3D whole-uterus MRI volumes in 153 pregnant women acquired at the Boston Children’s Hospital. We train another instance of our method on dHCP, a publicly available dataset of fetal brain MRI scans in 245 subjects [karolis2025]. Subject gestational age (GA) is 18-38 and 20-38 weeks in Research-Fetal and dHCP, respectively. All volumes have 2-3mm isotropic voxels. They are not interleaved with 2D diagnostic slices and do not contain spin history artifacts. We use training/validation/testing splits of 114/15/25 and 148/48/49 subjects for Research-Fetal and dHCP, respectively. dHCP consists of higher quality volumes, presenting a more challenging clinical domain generalization task compared to Research-Fetal. Additionally, dHCP volumes are neatly cropped around the fetal head. Therefore, segmentation networks trained on dHCP fail to accurately predict the brain mask in volumes where the field of view is larger and includes uterine regions outside the head and maternal tissues (Fig. 3). For this reason, we train our segmentation network only on Research-Fetal.
@@
We evaluate the methods on Clinical-Young, a dataset of routine clinical volumes in 60 younger fetuses (GA 18-23 weeks). This dataset presents two challenges: clinical volumes exhibit a substantial domain shift from high-quality research data, and underdeveloped anatomy complicates pose estimation [zhan2013]. The voxel size is 1.76-3.5mm, and volumes are not interleaved with diagnostic slices. We also test the models on Navigators, a dataset containing 47 time-series (1210 total volumes) of 3D MRI navigator volumes interleaved with 2D slices in 9 subjects (GA 26-36 weeks). Due to fast and interleaved acquisitions, these volumes have voxel size 4-6mm and contain real spin history artifacts. This dataset represents the largest domain gap from the research data, and volumes are representative of the intended clinical application, i.e, adaptive slice prescription. Both clinical datasets are acquired at the Boston Children’s Hospital.
@@ 5.2 Evaluation metrics
We assess performance with the geodesic rotation error $\arccos{(\frac{1}{2}[\text{trace}(R\hat{R}^{-1})-1])}$ and the average absolute distance (AAD) for voxels on the brain surface [greve2009] (see Appendix G for additional metrics).
@@ 5.3 Baseline methods
We evaluate two baseline methods that perform rigid registration to a template volume. FireANTs [jena2024] is a state-of-the-art method for fast, optimization-based 3D image registration. EquiTrack [billot2024] uses a SE(3)-equivariant network that learns matching landmarks to compute the optimal rigid transform from the input to the template via SVD [levinson2020]. To enable comparison with E(3)-Pose, we generate subject-specific template volumes using the GT pose. For accurate registration, we mask input volumes with the brain masks predicted by our segmentation network.
@@
We evaluate four baseline methods that estimate the canonical pose from only the input volume. Fetal-Align [hoffmann2021] implements landmark-based pose estimation, which segments the brain and eyes as landmarks and uses brain shape to resolve left-right ambiguity. For fair comparison on our datasets, we use our trained segmentation network, which we train to segment both the brain and eyes, to perform landmark detection. The remaining three baselines directly regress pose with CNNs. 3DPose-Net [salehi2019] and 6DRep [faghihpirayesh2023] regress the axis-angle vector and two basis directions of the object frame, respectively. The latter method applies Gram-Schmidt orthonormalization [bjorck1967] to the network output to obtain the final rotation. Registration-by-Regression (RbR) [gopinath2024] trains a U-Net to regress a deformation field to the object frame at the same resolution as the input volume. The final rotation is computed from the network output with SVD [horn_svd_1987]. We train all three networks on the research datasets and tune augmentation parameters for optimal performance. We evaluate rotation only and use our CoM translation estimation for all methods. Appendix F provides further details on the baseline methods.

### v5

@@ 4 Methods
Our goal is to estimate a 6-DoF canonical object pose $T$ from a 3D volume, with respect to the volume coordinate frame. We estimate the rigid pose $T\!=\!t\circ R$ by separating it into its rotation $R$ and translation $t$ components (Fig. 2).
@@ 4.1 Translation estimation
In rigid pose estimation, it is customary to place the origin of the object coordinate frame at the object center [75]. If $\psi:\mathbb{R}^{3}\rightarrow\{0,1\}$ is a binary object occupancy function, then $t$ can be computed from a set $X$ of sampled points (i.e., voxel coordinates): $t=\frac{\sum_{x\in X}x\psi(x)}{\sum_{x\in X}\psi(x)}.$ See Appendix C for our implementation of $\psi$ using a CNN for fetal brain segmentation from whole-uterus MRI volumes.
@@ 4.2 Rotation estimation
Let $G_{\text{symm}}\!=\!\{g_{s}\in\text{O(3)}\!:\!f\!\circ\!g_{s}^{-1}\!=\!f\}$ be the object symmetry group [62]. To handle object symmetries, we aim to parametrize $h(R)$ such that $\rho_{h}(g_{s})h(R)\!=\!h(R)$ for all $g_{s}\!\in\!G_{\text{symm}}$ and $R\!\in\!\text{SO(3)}$; and $h$ remains continuous over SO(3). Here, we consider the case $G_{\text{symm}}\!=\!\{g_{\leftrightarrow}\}$, where $g_{\leftrightarrow}$ denotes left-right reflection, defined by the axis of left-right symmetry $e_{\leftrightarrow}$.
Formally, we define $h(R)$ based on three orthonormal basis directions $\{e_{x},e_{y},e_{z}\}$ of the object frame (d=9d\!=\!9):
@@
where $\oplus$ denotes columnwise/block-diagonal concatenation for vectors/matrices, respectively. Specifically, we decompose $h(R)$ into three irreducible tensors: a pseudovector parametrizes the left-right direction, and two vectors parametrize two directions orthogonal to the left-right axis. Invariance under $G_{\text{symm}}$ is straightforward:
$\rho_{h}(g_{\leftrightarrow})h(R)$ $=\!(\det{M(g_{\leftrightarrow})})M(g_{\leftrightarrow})e_{x}\!\oplus\!M(g_{\leftrightarrow})e_{y}\!\oplus\!M(g_{\leftrightarrow})e_{z},$ (6)
@@
where the second equality follows from the fact that following reflection, vectors orthogonal to the reflection axis are unchanged, and those that are parallel become inverted. Since we use an E(3)-CNN to estimate $h(R)$, we ensure that $h(R)$ is equivariant under $\text{SE(3)}\!\subset\!\text{E(3)}$. $\rho_{h}$ is essentially the rotation matrix when group actions are restricted to SO(3), establishing the continuity of $h$ over SO(3).
@@
where $(e_{x},e_{y},e_{z})$ is the ground-truth (GT) orthonormal basis, $\theta_{k}\!=\!\arccos{(\hat{e}_{k}\cdot e_{k})}$ for $k\in\{x,y,z\}$, and $\beta$ is a weight hyperparameter. Our loss function is invariant under $G_{\text{symm}}$ (Appendix D). During inference, we obtain an orthonormal output matrix by applying SVD to the predicted basis [45]. We then choose between $\hat{e}_{x}$ and $\hat{e}_{x}$ in order to guarantee a positive determinant matrix, i.e., a proper rotation. This step is related to prior work that heuristically selects the output pose from a set of symmetrically plausible poses [56, 32, 81]. Here, the task is straightforward since we are choosing between a right-handed and left-handed rotation.
@@ 4.3 Implementation details
We augment training volumes with commonly used spatial (rigid transformations, scaling, flips) and intensity (bias field, gamma, Gaussian noise, low resolution) transforms [55, 4] in volumetric brain MRI. To enable our method to generalize to navigator volumes interleaved with diagnostic slices, we additionally simulate a spin history artifact from the preceding slice [50, 18], which appears as dark shading along the slice imaging plane (blue arrows in Fig. 3). We model this artifact with a Gaussian approximation of the slice profile $\tilde{f}(x)\!=\!f(x)(1-\mathcal{N}((x-c_{\text{slice}})^{T}n_{\text{slice}};0,\sigma^{2})$, where $f$ and $\tilde{f}$ are the original and augmented volumes, respectively; $c_{\text{slice}}$ and $n_{\text{slice}}$ are the center and normal to the slice imaging plane, sampled uniformly at random; and $\sigma^{2}$ is determined by slice thickness and the rate of spin decay [40] (see Appendix F for more details).
Table 1: Baseline comparisons. Mean $\pm$ standard deviation for rotation error () and average absolute error (AAD, $m$). Best score is in bold. * indicates statistical significance at p<0.05p\!<\!0.05 compared to E(3)-Pose (hierarchical permutation test for Navigators, pairwise Wilcoxon for all other datasets). FireANTs and Fetal-Align are not re-trained for rotation estimation; we report the performance statistics separately for these methods. FireANTs runtime is 1.8s, prohibiting real-time deployment. EquiTrack runtime is 0.7s, and the remaining methods run in 0.3s. See Tables 4-5 for subject-level statistics in Navigators.
@@
FireANTs [36] 10.6 $\pm$ 24.5* 5.7 $\mathbf{\pm}$ 9.3* 0.4 $\mathbf{\pm}$ 1.6* 0.6 $\mathbf{\pm}$ 1.0* 17.3 $\pm$ 40.6 4.1 $\pm$ 7.4 44.6 $\pm$ 55.6* 19.1 $\pm$ 20.3*
Fetal-Align [33] 5.0 $\mathbf{\pm}$ 3.0 3.0 $\mathbf{\pm}$ 1.7 8.5 $\pm$ 24.7* 3.5 $\pm$ 6.8* 10.3 $\pm$ 21.0* 3.2 $\pm$ 5.0* 56.2 $\pm$ 57.6* 24.1 $\pm$ 23.3*
@@
EquiTrack [3] 11.3 $\pm$ 13.1* 6.2 $\pm$ 5.9* 7.1 $\pm$ 5.0 3.6 $\pm$ 2.6 18.9 $\pm$ 29.0* 5.4 $\pm$ 5.7* 44.7 $\pm$ 48.7* 20.3 $\pm$ 18.3*
3DPose-Net [60] 19.8 $\pm$ 15.0* 10.5 $\pm$ 7.6* 24.0 $\pm$ 19.2* 10.9 $\pm$ 7.5* 39.2 $\pm$ 36.8* 11.2 $\pm$ 8.6* 65.0 $\pm$ 47.8* 27.7 $\pm$ 16.3*
6DRep [16] 9.6 $\pm$ 4.5* 5.3 $\pm$ 2.3* 9.8 $\pm$ 5.5* 4.9 $\pm$ 3.0* 14.1 $\pm$ 7.7* 4.6 $\pm$ 2.3* 38.8 $\pm$ 45.7* 17.2 $\pm$ 15.9*
RbR [25] 7.6 $\pm$ 3.9* 4.3 $\pm$ 2.2* 8.6 $\pm$ 5.0 4.3 $\pm$ 2.6 11.2 $\pm$ 4.6* 3.7 $\pm$ 1.4* 22.6 $\pm$ 24.1* 11.5 $\pm$ 9.9*
@@
EquiTrack [3] 29.5 $\pm$ 43.5* 14.1 $\pm$ 19.5* 13.6 $\pm$ 36.7* 5.4 $\pm$ 11.6* 48.3 $\pm$ 62.1* 11.6 $\pm$ 13.2* 59.0 $\pm$ 57.5* 25.7 $\pm$ 22.1*
3DPose-Net [60] 33.5 $\pm$ 26.6* 17.1 $\pm$ 11.6* 21.2 $\pm$ 27.4* 9.0 $\pm$ 8.9* 51.5 $\pm$ 37.8* 14.4 $\pm$ 8.9* 93.9 $\pm$ 55.8* 36.2 $\pm$ 16.7*
6DRep [16] 32.2 $\pm$ 26.0* 17.3 $\pm$ 14.2* 12.2 $\pm$ 6.1* 5.9 $\pm$ 3.1* 48.2 $\pm$ 50.8* 12.4 $\pm$ 10.8* 84.1 $\pm$ 51.2* 34.0 $\pm$ 17.3*
RbR [25] 15.4 $\pm$ 18.3* 8.3 $\pm$ 8.5* 8.0 $\pm$ 4.0* 3.9 $\pm$ 1.9 50.9 $\pm$ 48.4* 12.7 $\pm$ 9.9* 77.2 $\pm$ 53.6* 30.8 $\pm$ 17.3*
@@ 5.1 Data
We train two variants of E(3)-Pose on two different large, high-quality research datasets. We first evaluate on in-distribution held-out test data. Then, we evaluate on two clinical datasets to assess domain generalization and the potential for clinical translation. All volumes are annotated for ground-truth pose and segmentations (Appendix G).
@@
We train one instance of E(3)-Pose on Research-Fetal, a dataset of 3D whole-uterus MRI volumes in 153 pregnant volunteers acquired at the Boston Children’s Hospital. We train another instance of our method on dHCP, a publicly available dataset of fetal brain MRI scans in 245 volunteers [37]. Fetus gestational age (GA) is 18-38 and 20-38 weeks in Research-Fetal and dHCP, respectively. All volumes have 2-3mm isotropic voxels. They are not interleaved with diagnostic slices and do not contain spin history artifacts. We use training/validation/testing splits of 114/15/25 and 148/48/49 participants for Research-Fetal and dHCP, respectively. dHCP consists of higher quality volumes, presenting a more challenging clinical domain generalization task compared to Research-Fetal. Additionally, dHCP volumes are neatly cropped around the fetal head. Therefore, segmentation networks trained on dHCP fail to accurately predict the brain mask in volumes where the field of view is larger and includes uterine regions outside the head and maternal tissues (Fig. 3). For this reason, we train our segmentation network only on Research-Fetal.
@@
We evaluate the methods on Clinical-Young, a dataset of routine clinical volumes in 60 younger fetuses (GA 18-23 weeks). This dataset presents two challenges: clinical volumes exhibit a substantial domain shift from high-quality research data, and underdeveloped anatomy complicates pose estimation [79]. The voxel size is 1.76-3.5mm, and volumes are not interleaved with diagnostic slices. We also test the models on Navigators, a dataset containing 47 time-series (1210 total volumes) of 3D MRI navigator volumes interleaved with 2D slices in 9 volunteers (GA 26-36 weeks). Due to fast and interleaved acquisitions, these volumes have voxel size 4-6mm and contain real spin history artifacts. This dataset represents the largest domain gap from the research data, and volumes are representative of the intended clinical application, i.e, adaptive slice prescription. Both clinical datasets are acquired at the Boston Children’s Hospital.
@@ 5.2 Evaluation metrics
We assess performance with the geodesic rotation error $\arccos{(\frac{1}{2}[\text{trace}(R\hat{R}^{-1})-1])}$ and the average absolute distance (AAD) for voxels on the brain surface [28] (see Appendix I for additional metrics).
@@ 5.3 Baseline methods
We evaluate two baseline methods that perform rigid registration to a template volume. FireANTs [36] is a state-of-the-art method for fast, optimization-based 3D image registration. EquiTrack [3] uses a SE(3)-equivariant network that learns matching landmarks to compute the optimal rigid transform from the input to the template via SVD [45]. To enable comparison with E(3)-Pose, we generate subject-specific template volumes using the GT pose. For accurate registration, we mask input volumes with the brain masks predicted by our segmentation network.
@@
We evaluate four baseline methods that estimate the canonical pose from only the input volume. Fetal-Align [33] implements landmark-based pose estimation, which segments the brain and eyes as landmarks and uses brain shape to resolve left-right ambiguity. For fair comparison on our datasets, we use our trained segmentation network, which we train to segment both the brain and eyes, to perform landmark detection. The remaining three baselines directly regress pose with CNNs. 3DPose-Net [60] and 6DRep [16] regress the axis-angle vector and two basis directions of the object frame, respectively. The latter method applies Gram-Schmidt orthonormalization [5] to the network output to obtain the final rotation. Registration-by-Regression (RbR) [25] trains a U-Net to regress a deformation field of the object frame at the same resolution as the input volume. The final rotation is computed from the network output with SVD [34]. We train all three networks on the research datasets and tune augmentation parameters for optimal performance. To run full rigid pose estimation for these methods, which estimate rotation only, we use CoM-based translation estimation. Appendix H provides further details on the baseline methods.